40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jul 22 16:30:31 2013, Koji, Update, General, Vent preparation 
    Reply  Mon Jul 22 16:56:14 2013, Jenne, Update, General, Vent preparation - In progress 
       Reply  Mon Jul 22 17:03:25 2013, Jenne, Update, General, MC spot positions MCdecenter_22July2013.png
          Reply  Tue Jul 23 04:26:54 2013, Jenne, Update, General, MC spot positions MCdecenter_22July2013_lowpower.png
       Reply  Mon Jul 22 18:49:16 2013, Jenne, Update, General, Vent preparation - In progress 
       Reply  Mon Jul 22 21:23:04 2013, manasa, Update, General, Vent preparation - In progress 
          Reply  Tue Jul 23 03:10:57 2013, Koji, Update, General, Vent preparation - In progress OSEM_RECORD.pngOPLEVS.png
          Reply  Tue Jul 23 03:43:32 2013, manasa, Update, General, Vent preparation - In progress 
             Reply  Tue Jul 23 04:22:03 2013, Koji, Update, General, Vent preparation - In progress 
       Reply  Tue Jul 23 04:31:36 2013, Jenne, Update, General, Vent preparation - Ready for Vent 
          Reply  Tue Jul 23 08:21:31 2013, Steve, Update, VAC, Vent has started beforVent.pngstartVent.png
             Reply  Tue Jul 23 14:00:03 2013, Steve, Update, VAC, the vent is completed Atm.png5cyl.pngventSum1d.png
Message ID: 8901     Entry time: Tue Jul 23 04:22:03 2013     In reply to: 8899
Author: Koji 
Type: Update 
Category: General 
Subject: Vent preparation - In progress 

[Koji Jenne]

Low power MC locking

- Rotated HWP right after the laser

- Put a knife edge beam dump at the output of the PBS after the HWP.

- Replaced the PO mirror for the MC refl by an HR mirror.

- PMC:
Input offset from 0 to 0.29
Servo Gain from 10 to 30
=> Transmission 0.84 (1.2W at the MC input) to 0.069 (100mW)

- MC:

VCO Gain from 25 to 31
MC REFL: Unlocked 3.6 Locked 0.38-0.40

ELOG V3.1.3-